Simply each expression. Write your final answer in a+bi form. Rewrite the expression so that it is an addition problem.
(3-8i)-(6+2i)
I tried doing (3-8i)-(-6+2i) I don't know if it is right and just needed some help on how to rewrite it.
Thanks!

Answer by Alan3354(69443)

Simply each expression. Write your final answer in a+bi form. Rewrite the expression so that it is an addition problem.
(3-8i)-(6+2i)
--------------
3-6 = -3
-8i - 2i = -10i
--> -3 - 10i